# On the dynamics of leapfrogging vortex rings

### v1

@@ 2.3. Expansion of the Biot-Savart law
In relation with (2.5), there is an explicit formula of the Green kernel, namely it is given in [26, Section 2] that
@@
This function $F$ was intensively studied in the literature. We bring together in the next lemma some estimates we need, and we refer to [26, Lemma 2.1] and [23, Lemma 2.7] for more details.
@@
$\max\Bigg{(}\Big{|}F(s)+\frac{1}{2}\ln(s)\Big{|},\Big{|}F^{\prime}(s)+\frac{1}{2s}\Big{|},|s^{2}F^{\prime\prime}(s)|\Bigg{)}\leqslant C_{0}.$
@@
Let $K$ a compact subset of +${\mathbb{R}}^{2}_{+}$, then there exists $R_{0}>0$ such that for any $x,y\in K$ we have $\frac{|x-y|^{2}}{x_{r}y_{r}}\leqslant R_{0}$. Hence, by Lemma 2.4, one has for all $x,y\in K$
$\Bigg{|}F\left(\frac{|x-y|^{2}}{x_{r}y_{r}}\right)+\ln|x-y|\Bigg{|}=\Big{|}\ln(\sqrt{x_{r}y_{r}})+\mathcal{O}(1)\Big{|}\leqslant C_{K}.$
@@
$=-\frac{x-y}{|x-y|^{2}}+\mathcal{O}(1)+\mathcal{O}\left(\frac{1}{x_{r}}\right),$
@@
$\Bigg{|}\nabla_{x}F\left(\frac{|x-y|^{2}}{x_{r}y_{r}}\right)+\nabla_{x}\ln|x-y|\Bigg{|}\leqslant C_{K}$
for all $x,y\in K$. As we have the same estimates for $\nabla_{y}$, this ends the proof. ∎
This expansion of the Green’s function together with (2.8) allows one to derive a decomposition of the velocity field $u$. For further reference, we denote the most singular part of the stream function
@@
Let $K$ a compact subset of +${\mathbb{R}}^{2}_{+}$ and $\omega\in L^{\infty}({\mathbb{R}}_{+}^{2})$ with $\operatorname{supp}(\omega)\subset K$. There exists $C_{K}>0$ depending only on $K$ such that the velocity field $u$ defined by (2.8) satisfies
@@
where $\|u_{R}\|_{L^{\infty}(K)}\leqslant C_{K}\int y_{r}\omega(y)\mathrm{d}y$.
@@
$\nabla^{\perp}\mathcal{G}(x,y)=\frac{\sqrt{x_{r}y_{r}}}{2\pi}\frac{(x-y)^{\perp}}{|x-y|^{2}}-\frac{1}{4\pi}\sqrt{\frac{{y_{r}}}{{x_{r}}}}\ln|x-y|e_{z}+\nabla^{\perp}_{x}S(x,y),$
hence the decomposition of $u$ comes from (2.8) and (2.10). The desired estimate of $u_{R}$ follows the estimates of $S$ in Corollary 2.5. ∎
@@ 3. A vortex ring under the influence of the other rings
as soon as $\varepsilon\leqslant\varepsilon_{0}$ for $\varepsilon_{0}>0$ depending on $h_{0}$.
Let $T_{0}>0$ to be chosen later. Defining
(3.6) $T_{\varepsilon}:=\sup\Big{\{}t\in[0,T_{0}],\,\forall s\in[0,t],\ \forall i\in\{1,\ldots,N\},\ \operatorname{supp}\omega_{i}^{\varepsilon}(s,\cdot)\subset\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)\Big{\}}$
which verifies, by continuity of the trajectories, $T_{\varepsilon}>0$ for every $\varepsilon>0$. Note that in particular, up to choosing $\varepsilon_{0}$ small enough, we have
@@
Provided $h_{0}\leqslant 1$, there exists $\varepsilon_{0}>0$ depending only on $X^{*}$, $(Y_{i,0})$ and $M_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, for every $i\in\{1,\ldots,N\}$, every $x,y\in\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)$ and every $t\in[0,T_{\varepsilon}]$,
@@
In this respect, it is crucial to verify that all the estimates below are uniform with respect to $i$. Since the proofs will focus only on one patch with one exterior field, we will often drop the index $i$ for the sake of readability.
In the sequel, every constant $C,C_{0},\varepsilon_{0}$ may depend on $X^{*}$, $(Y_{i,0})$ and $M_{0}$ and we omit to mention this dependence. At the opposite, we will carefully track the dependence on $h_{0}$ and $T_{0}$. Finally, we consider from now on that $\gamma>0$, and the opposite sign will be reached simply by reversing the time.
@@ 4. Weak localization
In this section, we focus on the weak localization which corresponds to Item (ii) in Theorem 1.1. The key proposition here is the analog of [49, Lemma 2.1], also used in a different context in [28].
@@
Provided $T_{0}\leqslant 1/(26C^{\prime}_{F})$ and $h_{0}\leqslant\min(r^{*}/240,1/2)$ the following holds true: for any positive function $\varepsilon\mapsto\rho(\varepsilon)$ such that $\rho(\varepsilon)\sqrt{|\ln\varepsilon|}$ tends to zero as $\varepsilon\to 0$, there exists $\varepsilon_{\rho}>0$, depending only on $\rho$ and $h_{0}$, such that for any $\varepsilon\in(0,\varepsilon_{\rho}]$, every $i\in\{1,\ldots,N\}$ and all $t\in[0,T_{\varepsilon}]$ there is an open set $\Omega_{\rho,i}^{\varepsilon}(t)$ satisfying
@@
Even if the proof was originally inspired by [49], we need to add several new arguments due to the stronger singularity of our problem. Indeed, the vortex patches are here close to each other, with a distance of order $1/\sqrt{|\ln\varepsilon|}$. The strategy in this section is to first assume that weak localization holds up to some time $\widetilde{T}_{\varepsilon}$, which is less precise than the desired estimates. This will allow to obtain sharp energy estimates and to deduce a sharper weak localization property on $[0,\widetilde{T}_{\varepsilon}]$, which will imply that $\widetilde{T}_{\varepsilon}=T_{\varepsilon}$ and that all the estimates proved up to $\widetilde{T}_{\varepsilon}$ hold actually true up to $T_{\varepsilon}$. Among these estimates and as they will be used later, we state a lemma encompassed in this bootstrap argument, which will be established in the proof of Proposition 4.1. This lemma involves the center of vorticity of each blob, which was defined at relation 1.12.
@@
Provided $T_{0}\leqslant 1/(26C^{\prime}_{F})$ and $h_{0}\leqslant\min(r^{*}/240,1/2)$, there exists $\varepsilon_{0}$ depending on $h_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$,
@@
Provided $T_{0}\leqslant 1/(26C^{\prime}_{F})$ and $h_{0}\leqslant\min(r^{*}/240,1/2)$ the following holds true: there exists $\varepsilon_{0}>0$ depending on $h_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$ and $i\in\{1,\ldots,N\}$,
@@ 4.1. Weak localization bootstrapping time
(4.2) $\frac{1}{\gamma}\int_{{\mathbb{R}}^{2}_{+}\setminus\Omega_{i}^{\varepsilon}(t)}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}x<\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\quad\text{and}\quad\operatorname{diam}\Omega_{i}^{\varepsilon}(t)<\frac{1}{|\ln\varepsilon|}.$
@@
The different localization properties related to the definition of $T_{\varepsilon}$ (3.6) and $\widetilde{T}_{\varepsilon}$ are illustrated in Figure 1. As often in the rest of this paper, we drop during the proofs the index $i$ in the notation of $\omega_{i}^{\varepsilon}$ and $\Omega_{i}^{\varepsilon}$ since we focus on a single vortex ring in an exterior field.
$r$$z$$X^{\varepsilon}(0)$$X^{\varepsilon}(t)$$2h_{0}$$\frac{2d_{0}}{\sqrt{|\ln\varepsilon|}}$$t\leqslant T_{\varepsilon}$trong localization
@@
Figure 1. Weak and strong localization: on the left, the support of $\omega^{\varepsilon}$ remains within a disk of size $\mathcal{O}(1)$ for $t\leqslant T_{0}$, and within a strip of width $2d_{0}/\sqrt{|\ln\varepsilon}$ for $t\leqslant T_{\varepsilon}$. On the right, most of the mass is within a circle of diameter $1/|\ln\varepsilon|$ for $t\leqslant\widetilde{T}_{\varepsilon}$, while initially all the mass was within a circle of radius $\varepsilon$.
@@
Provided $h_{0}\leqslant 1$, there exists $\varepsilon_{0}>0$ depending on $h_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$, relation (4.1) holds true on the time interval $[0,\widetilde{T}_{\varepsilon}]$. In addition, we have on the same time interval that
@@
Let $x_{0}\in\Omega^{\varepsilon}(t)\cap\operatorname{supp}(\omega^{\varepsilon}(t,\cdot))$, hence $\Omega^{\varepsilon}(t)\subset B\Big{(}x_{0},\frac{1}{|\ln\varepsilon|}\Big{)}$. By definition of $b^{\varepsilon}$ (1.12) and by (3.3), we write
@@ 4.2. Expansion of the energy
Provided $h_{0}\leqslant 1/2$, there exists $\varepsilon_{0}>0$ depending on $h_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$, $t\in[0,\widetilde{T}_{\varepsilon}]$ and $x\in\operatorname{supp}\omega^{\varepsilon}_{i}(t)$,
@@
Since $t\leqslant T_{\varepsilon}$, then $\operatorname{diam}\big{(}\operatorname{supp}\omega^{\varepsilon}(t)\big{)}\leqslant 2h_{0}$, so provided that $h_{0}<1/2$, then
$\psi^{\varepsilon}(t,x)\leqslant 0\quad\text{if }x\in\operatorname{supp}\omega^{\varepsilon}(t).$
@@
$\hskip 28.45274pt+\frac{\sqrt{x_{r}}}{2\pi}\int\left(\sqrt{y_{r}}-\sqrt{b^{\varepsilon}_{r}(t)}\right)\ln|x-y|y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y$
$:=A_{1}+A_{2}.$
@@
(4.3) $-\int\ln|x-y|y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y$ $\leqslant-2\pi M\int_{0}^{\sqrt{\frac{\gamma}{\pi M}}}s\ln s\mathrm{d}s=-2\pi M\Big{(}\frac{1}{2}s^{2}\ln s-\frac{1}{2}s^{2}\Big{)}\Big{|}_{s=\sqrt{\frac{\gamma}{\pi M}}}$
@@
By the same computations as the ones in the proofs of [20, Lemma 4.10] and [31, Proposition 5.6], using the symmetry of $\mathcal{G}$ and the fact that $\omega^{\varepsilon}$ satisfies (3.2) in the sense of Proposition 2.1, we get
@@
$\hskip 28.45274pt+\frac{1}{2\pi|\ln\varepsilon|}\iint F^{\varepsilon}(t,x)\cdot e_{r}\sqrt{\frac{y_{r}}{x_{r}}}\ln|x-y|x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
$\hskip 56.9055pt+\frac{2}{|\ln\varepsilon|}\iint F^{\varepsilon}(t,x)\cdot\nabla_{x}S(x,y)x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
$:=A_{1}+A_{2}+A_{3}.$
@@
$|A_{1}|$ $\leqslant\frac{C_{F^{\prime}}}{2\pi}\iint\sqrt{x_{r}y_{r}}x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y\leqslant\frac{r^{\ast}C_{F^{\prime}}\gamma^{2}}{\pi}.$
@@
$\leqslant\frac{2C_{F}\gamma^{2}}{r^{*}2\pi\sqrt{|\ln\varepsilon|}}\left(|\ln\varepsilon||b_{r}^{\varepsilon}(0)-b_{r}^{\varepsilon}(t)|+\mathcal{O}(1)\right)$
@@
where we have used that $b^{\varepsilon}(0),b^{\varepsilon}(t)\in\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)$ (see (3.4) for the definition of this set), hence $|b_{r}^{\varepsilon}(0)-b_{r}^{\varepsilon}(t)|\leqslant 2d_{0}/\sqrt{|\ln\varepsilon|}$.
@@
Then, we obtain the following.
@@
Using (ii) in Proposition 2.1, we compute
@@
$\left|\mathcal{E}_{i}^{\varepsilon}(t)-\frac{\gamma^{2}}{2\pi}|\ln\varepsilon|b_{i,r}^{\varepsilon}(t)\right|\leqslant\frac{25C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|.$
@@
Gathering the results of Lemmas 4.8 and 4.9, we have that for every $t\leqslant\widetilde{T}_{\varepsilon}$,

### v2

@@ 2.3. Expansion of the Biot-Savart law
In relation with (2.7), there is an explicit formula of the Green kernel, namely it is given in [30, Section 2] that
@@
This function $F$ was intensively studied in the literature. We bring together in the next lemma some estimates we need, and we refer to [30, Lemma 2.1] and [27, Lemma 2.7] for more details.
@@
$\max\Bigg{(}\Big{|}F(s)+\frac{1}{2}\ln(s)\Big{|},\frac{1}{|\ln s|}\Big{|}F^{\prime}(s)+\frac{1}{2s}\Big{|},|s^{2}F^{\prime\prime}(s)|\Bigg{)}\leqslant C_{0}.$
@@
Let $\mathcal{K}$ a compact subset of +${\mathbb{R}}^{2}_{+}$, then there exists $R_{0}>0$ such that for any $x,y\in\mathcal{K}$ we have $\frac{|x-y|^{2}}{x_{r}y_{r}}\leqslant R_{0}$. Hence, by Lemma 2.4, one has for all $x,y\in\mathcal{K}$
$\Bigg{|}F\left(\frac{|x-y|^{2}}{x_{r}y_{r}}\right)+\ln|x-y|\Bigg{|}=\Big{|}\ln(\sqrt{x_{r}y_{r}})+\mathcal{O}(1)\Big{|}\leqslant C_{\mathcal{K}}.$
@@
$=-\frac{x-y}{|x-y|^{2}}+\mathcal{O}\left(\frac{1}{\sqrt{x_{r}y_{r}}}\right)+\mathcal{O}\left(\frac{1}{x_{r}}\right),$
@@
$\Bigg{|}\nabla_{x}F\left(\frac{|x-y|^{2}}{x_{r}y_{r}}\right)+\nabla_{x}\ln|x-y|\Bigg{|}\leqslant C_{\mathcal{K}}$
for all $x,y\in\mathcal{K}$. As we have the same estimates for $\nabla_{y}$, this ends the proof. ∎
This expansion of the Green’s function together with (2.11) allows one to derive a decomposition of the velocity field $u$. For further reference, we denote the most singular part of the stream function by
@@
Let $\mathcal{K}$ a compact subset of +${\mathbb{R}}^{2}_{+}$ and $\omega\in L^{\infty}({\mathbb{R}}_{+}^{2})$ with $\operatorname{supp}(\omega)\subset\mathcal{K}$. There exists $C_{\mathcal{K}}>0$ depending only on $\mathcal{K}$ such that the velocity field $u$ defined by (2.11) satisfies
@@
where $\|u_{R}\|_{L^{\infty}(\mathcal{K})}\leqslant C_{\mathcal{K}}\int y_{r}\omega(y)\mathrm{d}y$.
@@
$\nabla^{\perp}_{x}\mathcal{G}(x,y)=\frac{\sqrt{x_{r}y_{r}}}{2\pi}\frac{(x-y)^{\perp}}{|x-y|^{2}}-\frac{1}{4\pi}\sqrt{\frac{{y_{r}}}{{x_{r}}}}\ln|x-y|e_{z}+\nabla^{\perp}_{x}S(x,y),$
hence the decomposition of $u$ comes from (2.11) and (2.13). The desired estimate of $u_{R}$ follows the estimate of $S$ in Corollary 2.5. ∎
@@ 3. A vortex ring under the influence of the other rings
as soon as $\varepsilon\leqslant\varepsilon_{0}$ for $\varepsilon_{0}>0$ depending on $h_{0}$ and $d_{0}$.
Let $T_{0}>0$ to be chosen later. We define
(3.6) $T_{\varepsilon}:=\sup\Big{\{}t\in[0,T_{0}],\,\forall s\in[0,t],\ \forall i\in\{1,\ldots,N\},\ \operatorname{supp}\omega_{i}^{\varepsilon}(s,\cdot)\subset\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)\Big{\}},$
which verifies, by continuity of the trajectories, $T_{\varepsilon}>0$ for every $\varepsilon\in(0,\varepsilon_{0}]$. Note that in particular, up to choosing $\varepsilon_{0}$ small enough, we have
@@
Provided $h_{0}\leqslant 1$, there exists $\varepsilon_{0}>0$ depending only on $X^{*}$, $(Y_{i,0})$ and $M_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, every $i\in\{1,\ldots,N\}$, every $x,y\in\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)$ and every $t\in[0,T_{\varepsilon}]$,
@@
In this respect, it is crucial to verify that all the estimates below are uniform with respect to $i$. Since the proofs will focus only on one patch for a fixed $i\in\{1,\dots,N\}$, we will often drop the index $i$ for the sake of readability.
In the sequel, every constant $C,C_{0},\varepsilon_{0}$ may depend on $X^{*}$, $(Y_{i,0})$ and $M_{0}$ and we omit to mention this dependence. At the opposite, we will carefully track the dependence on $h_{0}$ and $T_{0}$. Finally, we consider from now on that $\gamma>0$, and the opposite sign will be reached simply by reversing the time. By (1.5), this means that we are focusing only on non-negative vortices.
@@ 4. Weak localization
[…]
@@
Given $T_{\varepsilon}>0$ as in (3.6), provided $T_{0}\leqslant 1/(24C^{\prime}_{F})$ and $h_{0}\leqslant\min(r^{*}/240,1/2)$ the following holds true: for any positive function $\varepsilon\mapsto\rho(\varepsilon)$ such that $\rho(\varepsilon)\sqrt{|\ln\varepsilon|}$ tends to zero as $\varepsilon\to 0$, there exists $\varepsilon_{\rho}>0$, depending only on $\rho$ and $h_{0}$, such that for any $\varepsilon\in(0,\varepsilon_{\rho}]$, every $i\in\{1,\ldots,N\}$ and all $t\in[0,T_{\varepsilon}]$ there is an open set $\Omega_{\rho,i}^{\varepsilon}(t)$ satisfying
@@
Even if the proof was originally inspired by [55] in the context of the two dimensional Euler equations, we need to add several new arguments due to the anelastic constraint and the stronger singularity of our problem. Indeed, the vortex patches are moving by the binormal flow, highlighted by the presence of $u_{L}^{\varepsilon}$ in Lemma 2.6, and are here close to each other, with a distance of order $1/\sqrt{|\ln\varepsilon|}$. The strategy in this section is to first assume that weak localization holds up to some time $\widetilde{T}_{\varepsilon}$, which is less precise than the desired estimates. This will allow to obtain sharp energy estimates and to deduce a sharper weak localization property on $[0,\widetilde{T}_{\varepsilon}]$, which will imply that $\widetilde{T}_{\varepsilon}=T_{\varepsilon}$ and that all the estimates proved up to $\widetilde{T}_{\varepsilon}$ hold actually true up to $T_{\varepsilon}$. Among these estimates and as they will be used later, we state a lemma encompassed in this bootstrap argument, which will be established in the proof of Proposition 4.1. This lemma involves the center of vorticity of each blob, which was defined at (1.6).
@@
Provided $T_{0}\leqslant 1/(24C^{\prime}_{F})$ and $h_{0}\leqslant\min(r^{*}/240,1/2)$, there exists $\varepsilon_{0}$ depending on $h_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$,
@@
Given $T_{\varepsilon}>0$ as in (3.6), provided $T_{0}\leqslant 1/(24C^{\prime}_{F})$ and $h_{0}\leqslant\min(r^{*}/240,1/2)$ the following holds true: there exists $\varepsilon_{0}>0$ depending on $h_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$ and $i\in\{1,\ldots,N\}$,
@@ 4.1. Weak localization bootstrapping time
(4.2) $\frac{1}{\gamma}\int_{{\mathbb{R}}^{2}_{+}\setminus\Omega_{i}^{\varepsilon}(t)}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}x\leqslant\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\quad\text{and}\quad\operatorname{diam}\Omega_{i}^{\varepsilon}(t)\leqslant\frac{1}{|\ln\varepsilon|}.$
@@
The different localization properties related to the definition of $T_{\varepsilon}$ (3.6) and $\widetilde{T}_{\varepsilon}$ are illustrated in Figure 2.
$z$$r$$X^{\varepsilon}(0)$$X^{\varepsilon}(t)$$2h_{0}$$\frac{2d_{0}}{\sqrt{|\ln\varepsilon|}}$$t\leqslant T_{\varepsilon}$trong localization
@@
Figure 2. Strong and weak localization: on the left, the support of $\omega^{\varepsilon}$ remains within a rectangle of height $2h_{0}$ and width $2d_{0}/\sqrt{|\ln\varepsilon}$ for $t\leqslant T_{\varepsilon}$. On the right, most of the mass is concentrated within a ball of diameter $1/|\ln\varepsilon|$ for $t\leqslant\widetilde{T}_{\varepsilon}$, while initially all the mass was contained in a ball of radius $\varepsilon$.
@@
Provided $h_{0}\leqslant 1$, there exists $\varepsilon_{0}>0$ depending on $h_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$, the inequality (4.1) holds true on the time interval $[0,\widetilde{T}_{\varepsilon}]$. In addition, it holds for all $[0,\widetilde{T}_{\varepsilon}]$ that
@@
Here and below, we frequently drop the index $i$ in the notation of $\omega_{i}^{\varepsilon}$ and $\Omega_{i}^{\varepsilon}$ during the proofs as they are performed for a fixed $i\in\{0,\ldots,N\}$ and we derive uniform estimates with respect to this parameter.
Let $x_{0}\in\Omega^{\varepsilon}(t)\cap\operatorname{supp}(\omega^{\varepsilon}(t,\cdot))$, hence $\Omega^{\varepsilon}(t)\subset B\Big{(}x_{0},\frac{1}{|\ln\varepsilon|}\Big{)}$. By definition of $b^{\varepsilon}$ (1.6) and by the conservation of the total mass of the vorticity (3.3), we write
@@ 4.2. Expansion of the energy
Provided $h_{0}\leqslant 1/2$, there exists $\varepsilon_{0}>0$ depending on $h_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$, $t\in[0,\widetilde{T}_{\varepsilon}]$ and $x\in\operatorname{supp}\omega^{\varepsilon}_{i}(t)$, it holds
@@
Since $t\leqslant T_{\varepsilon}$, we have $\operatorname{diam}\big{(}\operatorname{supp}\omega^{\varepsilon}(t)\big{)}\leqslant 2h_{0}$, so provided that $h_{0}<1/2$, then
$\psi^{\varepsilon}(t,x)\leqslant 0\quad\text{if }x\in\operatorname{supp}\omega^{\varepsilon}(t),$
where we recall that we are focusing only on non-negative vortices.
@@
$+\frac{\sqrt{x_{r}}}{2\pi}\int\left(\sqrt{y_{r}}-\sqrt{b^{\varepsilon}_{r}(t)}\right)\ln|x-y|y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y$
::=:= $A_{1}+A_{2}.$
@@
(4.3) $-\int\ln|x-y|y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y$ $\leqslant-2\pi M\int_{0}^{\sqrt{\frac{\gamma}{\pi M}}}s\ln s\mathrm{d}s=-2\pi M\Big{(}\frac{1}{2}s^{2}\ln s-\frac{1}{4}s^{2}\Big{)}\Big{|}_{s=\sqrt{\frac{\gamma}{\pi M}}}$
@@
By the same computations as the ones in the proofs of [23, Lemma 4.10] and [35, Proposition 5.6], using the symmetry of $\mathcal{G}$ and the fact that $\omega^{\varepsilon}$ satisfies (3.2) in the sense of Proposition 2.1, we get
@@
$+\frac{1}{2\pi|\ln\varepsilon|}\iint F^{\varepsilon}(t,x)\cdot e_{r}\sqrt{\frac{y_{r}}{x_{r}}}\ln|x-y|x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
$+\frac{2}{|\ln\varepsilon|}\iint F^{\varepsilon}(t,x)\cdot\nabla_{x}S(x,y)x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
::=:= $A_{1}+A_{2}+A_{3}.$
@@
$|A_{1}|$ $\leqslant\frac{C_{F^{\prime}}}{2\pi}\iint\sqrt{x_{r}y_{r}}x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y\leqslant\frac{r^{\ast}C_{F^{\prime}}\gamma^{2}}{\pi},$
@@
$\leqslant\frac{2C_{F}\gamma^{2}}{r^{*}2\pi\sqrt{|\ln\varepsilon|}}\Big{(}|\ln\varepsilon||b_{r}^{\varepsilon}(0)-b_{r}^{\varepsilon}(t)|+\mathcal{O}(1)\Big{)}$
@@
where we have used that $b^{\varepsilon}(0),b^{\varepsilon}(t)\in\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{0,r}}\left(d_{0},h_{0}\right)$ (see (3.4) for the definition of this set), hence $|b_{r}^{\varepsilon}(0)-b_{r}^{\varepsilon}(t)|\leqslant 2d_{0}/\sqrt{|\ln\varepsilon|}$.
@@
Next, we relate the radial velocity of the center of mass with the field generated by the other vortex rings.
@@
As for the computation of $\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}t}$ in the beginning of the proof of Proposition 4.8, we compute from the definition of $b^{\varepsilon}$ (1.6) and the equations verified by $\omega^{\varepsilon}$ (3.2) that
@@
$\left|\mathcal{E}_{i}^{\varepsilon}(t)-\frac{\gamma^{2}}{2\pi}|\ln\varepsilon|b_{i,r}^{\varepsilon}(t)\right|\leqslant\frac{(24T_{0}C^{\prime}_{F}+\frac{1}{4})\gamma^{2}}{\pi}h_{0}\ln|\ln\varepsilon|.$
@@
Gathering the results of Proposition 4.8 and Lemma 4.9, we have that for every $t\leqslant\widetilde{T}_{\varepsilon}$,